The problem says: without graphing tell Wethereach point is on the graph of the given line
12x+y=16;(1/2,3)

Found 2 solutions by macston, It is costly:
Answer by macston(5194)

.
The point is (1/2, 3) which means x=1/2 and y=3:
.
12x+y=16
.
Substitute for x and y:
12(1/2)+3=16
6+3=16
9=16 Nine does not equal sixteen.
.
The point given is not true for the equation,
thus it is not on the line.

Answer by It is costly(175)


given;(1/2,3)
12x+y=16
(12/1)(1/2)+3=16
6+3=16
9 does not=16 there for (1/2,3) is not on the given line.